# Unveiling Client Privacy Leakage from Public Dataset Usage in Federated Distillation

### v1

@@ 3.3. Membership Inference Attack
Co-op LiRA. Co-op LiRA is designed for scenarios where clients’ private datasets exhibit similar data distributions. In this case, the server can utilize the non-target clients’ local models as reference models to conduct MIA on the target client’s model. We notice that when target samples appear in other clients’ private datasets besides the target client, these clients’ private models cannot serve as reference models. However, in real-world FD deployments, clients are typically assumed to have mostly disjoint training datasets to benefit collaborative learning, thus making shared target samples rare. For Co-op LiRA, the server can conduct the proposed LDIA described in Section 3.2 to determine label distributions across all clients first. In the HFL setting, similar label distributions across clients indicate similar overall data distributions, resulting in comparable performance of local models. Then, the server can use other clients’ local models as reference models for LiRA directly based on the target client model. Algorithm 2 outlines the co-op LiRA process in PDA-FD: (1) During the communication phase, the server inserts the target sample into the selected public data subset to obtain the posterior probabilities from all clients’ models, including the target clients. (2) The server conducts LDIA on all clients. (3) The server calculates the KL-divergence between the label distributions of the target client and each remaining client, selecting clients as reference models when their KL-divergence falls below threshold $\beta$ (0.1). (4) The server performs LiRA hypothesis test by comparing reference models’ and target client’s posterior probabilities to determine target sample’s membership probability.
@@
(a) Reference model trained with $D_{ref}$.
(b) Reference model distilled by using $D_{ref}$.
@@
Distillation-based LiRA. To address the limitation of co-op LiRA, we proposed another extension of LiRA called distillation-based LiRA. It is challenging for the server to obtain a shadow dataset which has same data distribution with the target’s model’s training dataset to train a reference model; we instead choose to use knowledge distillation to generate a student model, which is then used as a reference model for the MIA. On the one hand, the generated student model should behave similarly to the target model, exhibiting close prediction scores on non-member samples. On the other hand, it produces different prediction scores on member data that differ from those from the teacher model, since the student model is trained through knowledge distillation without direct exposure to the training dataset, whereas the teacher model, trained directly on member samples, exhibits a degree of overfitting to these samples. Previous research on MIA (Jagielski et al., 2024) has demonstrated that the student model generated through knowledge distillation can potentially encode membership information of the teacher model. However, we find significant prediction discrepancies between student and teacher models when performing inference on the teacher model’s member samples. As shown in Figure 5(b), the prediction discrepancies between the teacher and the student model are significant for some members while remaining relatively small for non-members. This distinct pattern enables these distilled student models to serve as effective reference models for offline LiRA. These reference models allow attackers to identify members with high prediction discrepancies while maintaining a low False Positive Rate, thus achieving high True Positive Rate (TPR) at low False Positive Rate (FPR), which serves as a critical performance metric for successful membership inference attacks(Carlini et al., 2022; Watson et al., 2021). We validate this characteristic of knowledge distillation-based reference models in Section 4.3.
Algorithm 2 outlines the distillation-based LiRA process in PDA-FD: (1) During the communication phase, the server inserts the target samples into the public dataset. The server obtains the posterior probabilities predicted by the target client on all the data samples. These probability values could be used as soft labels to distill reference models. (2) To create multiple reference models, the server randomly samples the distillation dataset of distilled model from the public dataset except the target samples. Using the target client as the teacher model, the server distills multiple reference models. (3) The server then infers membership of the target samples by performing LiRA using the created reference models. This approach is more robust to heterogeneous data distributions among clients but comes at the cost of additional computational overhead for the distillation process.
@@ 4.1.1. Datasets
In our experiments, we utilize the following image datasets that are commonly used to test the performance of different FD frameworks: CIFAR-10(Krizhevsky et al., 2009), CINIC-10(Darlow et al., 2018), Fashion-MNIST(Xiao, 2017). Additionally, for the completeness of the experiments, we also use a tabular dataset:Purchase(DMDave et al., 2014). In our experiments, we partition each dataset into a 4:1 ratio for clients’ training sets $D_{train}$ and the public dataset $D_{pub}$. To align with the previous FD frameworks (Yang et al., 2022; Li and Wang, 2019), we also configure a scenario where there is a data distribution discrepancy between the public dataset and the clients’ private datasets. In this scenario, we use CIFAR-10 for client training and CIFAR-100 as the public dataset to simulate distribution shifts. Table 2 details the specific partitioning of datasets and the number of classes used in our experiments. In our MIA experiments, to ensure adequate private data for each client, we select an equal number of samples from the test dataset to serve as non-members.
Table 2. Datasets division.
Datasets number of classes $D_{train}$ $D_{pub}$ $D_{test}$
CIFAR-10 10 40000 10000 10000
CIFAR-10/CIFAR-100 10 40000 10000 10000
CINIC-10 10 72000 18000 90000
Fashion-MNIST 10 48000 12000 10000
Purchase 10 21589 5397 11565
@@
  • Mean $l1$-distance. The mean $l1$-distance represents the potential error between the inferred label distribution and the ground truth label distribution across all classes:
    (13) $Dis_{mean-l1}(\hat{p},p)=\frac{1}{M}\sum_{m=1}^{M}\mid\hat{p}_{m}-p_{m}\mid$
    A smaller mean $l1$-distance indicates a higher average accuracy of the LDIA results.
@@
To comprehensively evaluate the privacy leakage in the PDA-FD setting, we evaluate three different PDA-FD frameworks in our experiments: FedMD (Li and Wang, 2019), DS-FL (Itahara et al., 2021), and Cronus (Chang et al., 2019). Each of these FD frameworks employs a distinct approach to enhance the robustness of the FD algorithms. As mentioned in Section 2.1, they behave differently during the local updates phase and use different aggregation algorithms during the communication phase. In order to optimize the performance of all the PDA-FD frameworks, we should carefully select the number of training epochs in each round. Following the approach suggested by Li et al. in FedMD (Li and Wang, 2019), we initially train the local models to convergence on the private datasets before transitioning to the shorter update cycles during the distillation phase. Specifically, in the first round of local updates, each client performs 20 epochs of training. In the subsequent rounds, this is reduced to 5 epochs of training for each client. The knowledge distillation phase consists of 10 epochs of training for each client. To reduce communication cost, we randomly select 5000 data samples from the public dataset during each communication phase for the CIFAR-10, CIFAR10/CIFAR100, Fashion-MNIST and Purchase datasets, aligned with previous PDA-FD studies (Li and Wang, 2019). The CINIC-10 dataset, given its difficulty level, uses a set of 10000 randomly selected samples to ensure adequate knowledge transfer. Table 3 presents the performance of the three PDA-FD frameworks across various Dirichlet distributions and four distinct datasets.
@@ 4.2. Experiment Results of LDIA
Main Result. Table 4 presents the performance of the proposed LDIA on five different datasets across three PDA-FD frameworks. The results demonstrate the server’s capability to launch effective LDIA against clients across these datasets, significantly outperforming the random guess baseline on all three key metrics. For instance, for the DS-FL framework on the CIFAR-10 dataset with $\alpha$=1, our proposed LDIA achieves an average Chebyshev distance of 0.10, an average mean l1-distance of 0.03, and an average KL-divergence of 0.10 across all clients. In contrast, the random guess baseline yields substantially higher values: 0.20, 0.08, and 0.66 for the respective metrics. This significant improvement underscores the efficacy of our LDIA in accurately inferring clients’ label distributions. Figure 6 provides a visual representation of the LDIA results for the DS-FL server on the CIFAR-10 dataset, offering a clearer illustration of the experiment results. We can see from the figure that for the labels whose inferred proportions deviate from the ground truth values, the relative rankings of label frequencies are consistently preserved. This observation highlights the robustness of the proposed LDIA in capturing the essential structure of label distributions.
@@
Datasets Setting KL divergence Chebyshev distance Mean $l1$-distance
FedMD DS-FL Cronus Random FedMD DS-FL Cronus Random FedMD DS-FL Cronus Random
CIFAR10 $\alpha$=10 0.02 0.01 0.01 0.42 0.04 0.03 0.02 0.13 0.01 0.01 0.01 0.05
$\alpha$=1 0.17 0.10 0.08 0.66 0.14 0.11 0.10 0.20 0.04 0.03 0.03 0.08
$\alpha$=0.1 0.15 0.11 0.07 1.93 0.18 0.16 0.14 0.57 0.04 0.03 0.03 0.15
CIFAR10 $\alpha$=10 0.07 0.05 0.06 0.40 0.07 0.06 0.06 0.12 0.02 0.02 0.02 0.05
/CIFAR100 $\alpha$=1 0.11 0.10 0.10 0.59 0.10 0.11 0.11 0.22 0.03 0.03 0.03 0.08
$\alpha$=0.1 0.11 0.10 0.08 1.59 0.15 0.13 0.14 0.51 0.03 0.03 0.03 0.14
CINIC10 $\alpha$=10 0.01 0.01 0.01 0.64 0.02 0.02 0.04 0.15 0.01 0.01 0.01 0.07
$\alpha$=1 0.06 0.05 0.05 0.57 0.11 0.11 0.10 0.21 0.02 0.02 0.02 0.08
$\alpha$=0.1 0.09 0.08 0.01 1.99 0.14 0.14 0.04 0.56 0.03 0.03 0.01 0.15
Fashion $\alpha$=10 0.03 0.02 0.02 0.32 0.04 0.04 0.04 0.12 0.02 0.01 0.01 0.06
-MNIST $\alpha$=1 0.14 0.12 0.12 0.55 0.10 0.09 0.09 0.15 0.04 0.03 0.03 0.06
$\alpha$=0.1 0.21 0.05 0.06 1.54 0.20 0.09 0.11 0.45 0.04 0.02 0.02 0.13
Purchase $\alpha$=10 0.08 0.03 0.03 0.47 0.06 0.05 0.05 0.13 0.02 0.02 0.02 0.06
$\alpha$=1 0.27 0.14 0.15 0.68 0.13 0.10 0.10 0.18 0.06 0.04 0.04 0.09
$\alpha$=0.1 0.64 0.14 0.14 2.11 0.32 0.15 0.17 0.52 0.10 0.03 0.03 0.15
@@
Different Data Distributions. Our experiments reveal a notable relationship between the effectiveness of the proposed LDIA and the uniformity of clients’ label distributions. Specifically, the LDIA demonstrates lower KL-divergence, Chebyshev distance, and mean $l1$-distance as the clients’ label distributions become more uniform. This trend is clearly illustrated in our experiments using the CIFAR-10 dataset within the DSFL framework. As $\alpha$ increases, indicating a more uniform label distribution across clients, the LDIA achieves better performance. Conversely, as $\alpha$ decreases, indicating a more skewed distribution, we see an increase in the three key metrics. Nonetheless, the attack remains effective despite the reduced accuracy.
@@
We aggregate the server’s attack results across all clients for each round, to represent the overall LDIA performance over time. The results reveal that the server successfully executes LDIA on clients in every round. Notably, the LDIA performance remains relatively stable as the number of collaborative training rounds increases, showing neither significant improvement nor decline. This consistency can be attributed to the local updates phase preceding communication in each collaborative training round within PDA-FD frameworks. While this phase enhances knowledge transfer among clients, it simultaneously increases the degree of overfitting of each client’s local model to their private data. This dual effect contributes to the observed stability in LDIA performance over multiple rounds. Despite this stability, we recommend using the averaged LDIA results over multiple rounds as the final outcome to enhance the robustness of the results. Such an approach mitigates potential fluctuations and provides a more reliable measure of the server’s LDIA capabilities.
4.3. Experiment Results of MIA
In our MIA experiments, we evaluate the effectiveness of the proposed attack against 10 clients during the communication phase across three PDA-FD frameworks. These experiments use three different Dirichlet distributions and four distinct datasets. This comprehensive setup allows for a thorough assessment of MIA vulnerability under various data distribution scenarios and PDA-FD frameworks. Previous FD MIA studies (Wang et al., 2024; Liu and Dong, 2023) assume that attackers have access to shadow datasets matching the target model’s training data distribution. However, we find these assumptions too restrictive and unrepresentative of real-world scenarios. Therefore, our attack methods do not rely on such assumptions, so we do not use these studies as baselines for comparison.
@@ 5.1. Public Dataset Size
5.2. Number of Epochs in Local Updates Phase
Number of Epochs MIA (TPR at 1%FPR) LDIA (KL divergence)
2 8.10% 0.15
4 14.64% 0.10
6 15.43% 0.09

### v2

@@ 3.3.1. Co-op LiRA
Algorithm 2 outlines the co-op LiRA process in PDA-FD: (1) When the server prepares the public dataset, it incorporates a set of potential target samples into the public dataset. During the communication phase, the server selects subsets containing these target samples for knowledge transfer, allowing it to obtain posterior probabilities of the subset of public dataset from all clients’ models, including the target client. (2) The server conducts LDIA on all clients. (3) The server calculates the KL-divergence between the label distributions of the target client and each remaining client, selecting clients as reference models when their KL-divergence falls below threshold $\beta$ (0.1). (4) The server performs LiRA hypothesis test by comparing reference models’ and target client’s posterior probabilities on the target sample to determine target sample’s membership probability.
@@
(a) Reference model trained with $D_{ref}$
(b) Reference model distilled by using $D_{ref}$
@@
3.3.2. Distillation-based LiRA
To address the limitation of co-op LiRA, we proposed another extension of LiRA called distillation-based LiRA. It is challenging for the server to obtain a shadow dataset which has same data distribution with the target’s model’s training dataset to train a reference model; we instead choose to use knowledge distillation to generate a student model, which is then used as a reference model for the MIA. On the one hand, the generated student model should behave similarly to the target model, exhibiting close prediction scores on non-member samples. On the other hand, it produces different prediction scores on member data that differ from those from the teacher model, since the student model is trained through knowledge distillation without direct exposure to the training dataset, whereas the teacher model, trained directly on member samples, exhibits a degree of overfitting to these samples. Previous research on MIA (Jagielski et al., 2024) has demonstrated that the student model generated through knowledge distillation can potentially encode membership information of the teacher model. However, we find significant prediction discrepancies between student and teacher models when performing inference on the teacher model’s member samples. As shown in Figure 6(b), the prediction discrepancies between the teacher and the student model are significant for some members while remaining relatively small for non-members. This distinct pattern enables these distilled student models to serve as effective reference models for offline LiRA. These reference models allow attackers to identify members with high prediction discrepancies while maintaining a low False Positive Rate, thus achieving high True Positive Rate (TPR) at low False Positive Rate (FPR), which serves as a critical performance metric for successful membership inference attacks(Carlini et al., 2022; Watson et al., 2021). We validate this characteristic of knowledge distillation-based reference models in Section 4.3.
Algorithm 2 outlines the distillation-based LiRA process in PDA-FD: (1) When the server prepares the public dataset, it incorporates a set of potential target samples into the public dataset. During the communication phase, the server selects subsets containing the target samples for knowledge transfer. The server obtains the posterior probabilities predicted by the target client on all the data samples. These probability values could be used as soft labels to distill reference models. (2) To create multiple reference models, the server randomly samples the distillation dataset of distilled model from the public dataset except the target samples. Using the target client as the teacher model, the server distills multiple reference models. (3) The server then infers membership of the target samples by performing LiRA using the created reference models. This approach is more robust to heterogeneous data distributions among clients but comes at the cost of additional computational overhead for the distillation process.
@@ 4.1.1. Datasets
In our experiments, we utilize the following image datasets that are commonly used to test the performance of different FD frameworks: CIFAR-10(Krizhevsky et al., 2009), CINIC-10(Darlow et al., 2018), Fashion-MNIST(Xiao, 2017). Additionally, for the completeness of the experiments, we also use a tabular dataset:Purchase(DMDave et al., 2014). In our experiments, we partition each dataset into a 4:1 ratio for clients’ training sets $D_{train}$ and the public dataset $D_{pub}$. To align with the previous FD frameworks (Yang et al., 2022; Li and Wang, 2019), we also configure a scenario where there is a data distribution discrepancy between the public dataset and the clients’ private datasets. In this scenario, we use CIFAR-10 for client training and CIFAR-100 as the public dataset to simulate distribution shifts. The details of the specific partitioning of datasets and the number of classes used in our experiments can be found in Appendix A.1. In our MIA experiments, to ensure adequate private data for each client, we select an equal number of samples from the test dataset to serve as non-members.
@@
To comprehensively evaluate the privacy leakage in the PDA-FD setting, we evaluate three different PDA-FD frameworks in our experiments: FedMD (Li and Wang, 2019), DS-FL (Itahara et al., 2021), and Cronus (Chang et al., 2019). Each of these FD frameworks employs a distinct approach to enhance the robustness of the FD algorithms. As mentioned in Section 2.1, they behave differently during the local updates phase and use different aggregation algorithms during the communication phase. In order to optimize the performance of all the PDA-FD frameworks, we should carefully select the number of training epochs in each round. Following the approach suggested by Li et al. in FedMD (Li and Wang, 2019), we initially train the local models to convergence on the public and private datasets before transitioning to the shorter update cycles during the distillation phase. Specifically, in the first round of local updates, each client performs 20 epochs of training. In the subsequent rounds, this is reduced to 5 epochs of training for each client. The knowledge distillation phase consists of 10 epochs of training for each client. To reduce communication cost, we randomly select 5000 data samples from the public dataset during each communication phase for the CIFAR-10, CIFAR10/CIFAR100, Fashion-MNIST and Purchase datasets, aligned with previous PDA-FD studies (Li and Wang, 2019). The CINIC-10 dataset, given its difficulty level, uses a set of 10000 randomly selected samples to ensure adequate knowledge transfer. Table 2 presents the performance of the three PDA-FD frameworks across various Dirichlet distributions and five distinct dataset settings.
@@ 4.2. Experiment Results of LDIA
Main Result. Table 3 presents the performance of the proposed LDIA on five different datasets across three PDA-FD frameworks. The results demonstrate the server’s capability to launch effective LDIA against clients across these datasets, significantly outperforming the random guess baseline on all three key metrics. For instance, for the DS-FL framework on the CIFAR-10 dataset with $\alpha$=1, our proposed LDIA achieves an average Chebyshev distance of 0.11 and an average KL-divergence of 0.10 across all clients. In contrast, the random guess baseline yields substantially higher values: 0.20 and 0.66 for the respective metrics. This significant improvement underscores the efficacy of our LDIA in accurately inferring clients’ label distributions. Figure 7 provides a visual representation of the LDIA results for the DS-FL server on the CIFAR-10 dataset, offering a clearer illustration of the experiment results. We can see from the figure that for the labels whose inferred proportions deviate from the ground truth values, the relative rankings of label frequencies are consistently preserved. This observation highlights the robustness of the proposed LDIA in capturing the essential structure of label distributions.
@@
Datasets Setting KL divergence Chebyshev distance
FedMD DS-FL Cronus Random FedMD DS-FL Cronus Random
CIFAR10 $\alpha$=10 0.02 0.01 0.01 0.42 0.04 0.03 0.02 0.13
$\alpha$=1 0.17 0.10 0.08 0.66 0.14 0.11 0.10 0.20
$\alpha$=0.1 0.15 0.11 0.07 1.93 0.18 0.16 0.14 0.57
CIFAR10 $\alpha$=10 0.07 0.05 0.06 0.40 0.07 0.06 0.06 0.12
/CIFAR100 $\alpha$=1 0.11 0.10 0.10 0.59 0.10 0.11 0.11 0.22
$\alpha$=0.1 0.11 0.10 0.08 1.59 0.15 0.13 0.14 0.51
CINIC10 $\alpha$=10 0.01 0.01 0.01 0.64 0.02 0.02 0.04 0.15
$\alpha$=1 0.06 0.05 0.05 0.57 0.11 0.11 0.10 0.21
$\alpha$=0.1 0.09 0.08 0.01 1.99 0.14 0.14 0.04 0.56
Fashion $\alpha$=10 0.03 0.02 0.02 0.32 0.04 0.04 0.04 0.12
-MNIST $\alpha$=1 0.14 0.12 0.12 0.55 0.10 0.09 0.09 0.15
$\alpha$=0.1 0.21 0.05 0.06 1.54 0.20 0.09 0.11 0.45
Purchase $\alpha$=10 0.08 0.03 0.03 0.47 0.06 0.05 0.05 0.13
$\alpha$=1 0.27 0.14 0.15 0.68 0.13 0.10 0.10 0.18
$\alpha$=0.1 0.64 0.14 0.14 2.11 0.32 0.15 0.17 0.52
@@
Different Data Distributions. Our experiments reveal a notable relationship between the effectiveness of the proposed LDIA and the uniformity of clients’ label distributions. Specifically, the LDIA demonstrates lower KL-divergence, Chebyshev distance, and mean $l1$-distance as the clients’ label distributions become more uniform. This trend is clearly illustrated in our experiments using the CIFAR-10 dataset within the DS-FL framework. As $\alpha$ increases, indicating a more uniform label distribution across clients, the LDIA achieves better performance. Conversely, as $\alpha$ decreases, indicating a more skewed distribution, we see an increase in the three key metrics. Nonetheless, the attack remains effective despite the reduced accuracy.
@@
We aggregate the server’s attack results across all clients for each round, to represent the overall LDIA performance over time. The results reveal that the server successfully executes LDIA on clients in every round. Notably, the LDIA performance remains relatively stable as the number of collaborative training rounds increases, showing neither significant improvement nor decline. This consistency can be attributed to the local updates phase preceding communication in each collaborative training round within PDA-FD frameworks. While this phase enhances knowledge transfer among clients, it simultaneously increases the degree of overfitting of each client’s local model to their private data. This dual effect contributes to the observed stability in LDIA performance over multiple rounds.
4.3. Experiment Results of MIAs
In our MIA experiments, we evaluate the effectiveness of the proposed attack against 10 clients during the communication phase across three PDA-FD frameworks. Furthermore, we compare with existing attacks, including FD-leaks (Yang et al., 2022), MIA-FedDL (Liu and Dong, 2023), GradDiff (Wang et al., 2024), and Centralize-LiRA (Carlini et al., 2022), that serve as our baselines.
@@ 4.3. Experiment Results of MIAs
Comparison with Baselines. Our implementations of MIA-FedDL (Liu and Dong, 2023) and GradDiff (Wang et al., 2024) follow their proposed threat model settings, where the attacker can obtain a shadow dataset of data distribution consistent with the target model’s training dataset to train the shadow model. For the implementations of FD-leaks (Yang et al., 2022) and MIA-FedDL, we follow the same baseline experiment setup as described in GradDiff (Wang et al., 2024) with one modification: we reposition the attacker from the client side to the server side. For Centralize LiRA (Carlini et al., 2022), we chose to implement a more realistic threat model where the server directly performs LiRA on the target client model, with shadow models obtained solely through direct training on the already available public dataset. In our experiments, we chose to conduct experiments under the FedMD framework, with both the private dataset and the public dataset being CIFAR10. The parameter $\alpha$ of the Dirichlet distribution is set to 1. The results are shown in Table 6.
Table 6. Comparison with Baselines in FedMD(CIFAR10, $\alpha$=1).
Methods TPR at 1% FPR AUC Accuracy
FD-leaks (Yang et al., 2022) 0.00% 0.552 54.34%
MIA-FedDL (Liu and Dong, 2023) 1.75% 0.643 59.31%
GradDiff (Wang et al., 2024) 3.04% 0.695 65.17%
Centralize LiRA (Carlini et al., 2022) 6.46% 0.701 62.40%
Co-op LiRA* 7.21% 0.711 64.48%
Distillation-based LiRA* 17.62% 0.730 66.93%
    Our attack
From the table, we can see that our proposed Co-op LiRA and Distillation-based LiRA exhibit stronger attack performance, especially on TPR at low FPR region, one of the most important metrics for measuring MIA performance (Carlini et al., 2022). FD-leaks shows limited attack performance due to low-quality shadow models. As for MIA-FedDL and GradDiff, even under stronger and overly idealistic threat model conditions, they still demonstrate weaker attack results compared to LiRA attacks because they rely on classifier-based MIA techniques instead of incorporating state-of-the-art MIA methods. Meanwhile, our Co-op LiRA and Distillation-based LiRA outperform Centralized LiRA, further validating that our approaches for shadow model selection and training enable attackers to acquire higher-quality shadow models in realistic FD scenarios with constrained threat model assumptions, resulting in more effective attack results.
@@ 5.2. Number of Reference Models
5.3. Public-Private Data Distribution Shift
We also evaluate the impact of varying degrees of data distribution shift between the public and private datasets on both our attacks and the FD training process. We conduct our experiments following the FedMD federated distillation framework, where we set the Dirichlet distribution to $\alpha$ = 1 and the private dataset to CIFAR10. Then, we compare the attack efficacy of using different public datasets, including CIFAR10, CIFAR100, TinyImageNet (Le and Yang, 2015), and SVHN (Netzer et al., 2011). Table 8 shows the performance of FedMD and our attacks, LDIA and distillation-based LiRA, in different scenarios.
Table 8. Impact of Public-Private Data Distribution Shift.
Public Dataset MIA (TPR at 1%FPR) LDIA (KL divergence) FedMD Accuracy
CIFAR10 17.62% 0.17 75.38%
CIFAR100 9.47% 0.11 68.41%
TinyImageNet 7.83% 0.11 60.01%
SVHN 2.72% 0.18 47.12%
We can see that as the data distribution shift between public and private datasets increases, both attacks (LDIA and Distillation-based LiRA) become less effective. For LDIA, larger distribution shifts make it harder for the inference dataset $D_{inf}$ in LDIA to reflect the target model’s label sensitivity. For Distillation-based LiRA, increasing distribution shifts make it more difficult for shadow models to mimic target model performance by using knowledge distillation, reducing the attack effectiveness. Notably, we also find that larger distribution shifts simultaneously hinder the PDA-FD process itself, as effective knowledge distillation among clients becomes more difficult, ultimately degrading the overall federated learning performance.